OpenStudy (anonymous):

solve the system by the substituion method y=-5x-7 5x-2y=-1

OpenStudy (anonymous):

from the first equation, put -5x-7 in place of y

OpenStudy (anonymous):

If \(y=-5x-7\), then \[5x-2y=-1\]\[\Leftrightarrow5x-2(-5x-7)=-1\]Solve for x!

OpenStudy (anonymous):

5x-2(-5x-7)=-1 15x=-15 x=-1 y=-2

OpenStudy (anonymous):

i get how you got x but how is y-2

OpenStudy (anonymous):

it is give, y=-5x-7 put the value of x in this equation, u will get y

OpenStudy (anonymous):

it is given, sorry :)

OpenStudy (anonymous):

thanks its was a dumb question to obvious